### v1

# On the Robustness of Language Models for Tabular Question Answering

Abstract
Large Language Models (LLMs), originally shown to ace various text comprehension tasks have also remarkably been shown to tackle table comprehension tasks without specific training. While previous research has explored LLM capabilities with tabular dataset tasks, our study assesses the influence of in-context learning, model scale, instruction tuning, and domain biases on Tabular Question Answering (TQA). We evaluate the robustness of LLMs on Wikipedia-based WTQ and financial report-based TAT-QA TQA datasets, focusing on their ability to robustly interpret tabular data under various augmentations and perturbations. Our findings indicate that instructions significantly enhance performance, with recent models like Llama3 exhibiting greater robustness over earlier versions. However, data contamination and practical reliability issues persist, especially with WTQ. We highlight the need for improved methodologies, including structure-aware self-attention mechanisms and better handling of domain-specific tabular data, to develop more reliable LLMs for table comprehension.
On the Robustness of Language Models for Tabular Question Answering
[…]
Figure 1: Example of different possible table augmentation methods where each row operation represents original table augmentation, transpose-based table augmentation, and value-based perturbation. The value-based perturbation only affects the cell content associated with the answer value of the given question.
1 Introduction
LLMs, primarily trained on unstructured data, have performed fairly well with structured data like tabular datasets. Tasks that rely on tabular structures require the ability to process and understand data with structural awareness, even when the input is provided in a sequential linear format. Exploring such capability—particularly how LLMs apprehend and interpret structured information—can be effectively analyzed by studying robustness measures in table comprehension. This approach offers insights into the models’ ability to navigate and interpret the inherent relationships within structured datasets Borisov et al. (2023).
Tabular Question Answering (TQA) is a popular task on tables, which involves answering a query based on the given structured data Herzig et al. (2020); Fang et al. (2024). TQA task highlights the emergent properties of LLMs, demonstrating their capability to comprehend structured data through in-context learning, instruction adherence, and multi-step reasoning Fang et al. (2024); Wei et al. (2022); Wang et al. (2024); Jiang et al. (2023); Dong et al. (2024); Liu et al. (2024). This capability can be quantitatively evaluated by measuring the robustness of the model’s performance on tabular datasets. Wang et al. (2022) recently showed the vulnerability of LLMs to structural augmentation of the input table and introduced a novel methodology to leverage structure-aware self-attention to more effectively manage table transformations, thereby improving the performance of natural language generation from tables. Similarly, Zhao et al. (2023) identified the vulnerabilities of recent TQA LLMs, when faced with human-crafted adversarial perturbations. Liu et al. (2023) discusses inherent limitations within LLMs for tabular data interpretation, arguing that these models struggle with structural perturbations and advocate for an integrated approach that combines textual and symbolic reasoning. Chen (2023) find the performance of few-shot reasoning with LLMs for TQA show strong complex reasoning abilities over table structures without table-specific pre-training.
Although these studies showcase a critical need for enhanced methodologies in handling tabular data across various applications, they have yet to fully explore how the performance of robustness with the inherent ability of LLMs varies with different factors. Notably, there has been little comparison of robustness as a function of (1) in-context learning, (2) the scale and instruction-tuning of LLMs, and (3) the biases inherent within existing evaluation benchmarks for robust table-to-text generation. These aspects are crucial, as the scale of LLMs can influence their ability to generalize from data, while instruction-tuning enhances in-context learning ability, affecting how well they adapt to table comprehension without explicit re-training. Additionally, biases in training data can skew the evaluation of their comprehension ability. Hence, understanding these dimensions is essential for developing more reliable and fair LLMs in processing tabular data. To evaluate these, Wikipedia-based datasets are the most popular benchmark datasets for TQA. We experiment with WTQ dataset, a Wikipedia-based tabular question-answering dataset Pasupat and Liang (2015) and TAT-QA Zhu et al. (2021), a tabular question-answering dataset derived from financial reports. A general and specialized domain ensures reliable benchmarking of the robustness and generalizability of existing LLMs.
2 Various Perturbation Categories
Each perturbation is designed to manipulate table structure or content while preserving the inherent relational meaning of the table, to measure robustness to table comprehension.
2.1 Structural Perturbation (SP)
SP involves rearranging the columns and rows of the table to generate new examples. This ensures flexibility in understanding tabular data without distorting the semantics of the table. SP involves column swap, row swap, transpose, transpose column swap, and transpose row swap.
2.2 Value Perturbation (VP)
VP focuses on modifying the actual data values within tables, ensuring that the model is faithful to the table. We explore two types of VP:
Data Type Preserving Perturbation (DVP) DVP involves altering the answer to the question and, respectively, the cell values within the table while maintaining their original data types. For instance, in Fig 7, given a question, ‘‘What was the first venue for the Asian games?’’, we modify the correct answer, ‘‘Bangkok, Thailand", to ‘‘Beijing’’. We use such counterfactual entities to test faithfulness of the LLM to the tabular data. 111 We filter out the subset of data points where the table does not contain the answer, e.g., ‘‘How many people stayed at least 3 years in office?’’. We use an automated LLM-based counterfactual answer generation by prompting GPT-3.5, which ensures type-correctness of the altered answer.  [^2] The full prompts are present in the Appendix.
Random Value Perturbation (RVP) RVP (an example is shown in Fig. 8) relaxes DVP where instead of a counterfactual entity we have a fixed string, e.g., ‘‘r@nD0m v@1u3’’.
2.2.1 Null Value Perturbation (NVP)
NVP removes the correct answer from the table completely. Evaluating the performance on NVP highlights the influence of Wikipedia content on solving the WTQ table question-answering task.
2.2.2 No Table (NT)
To understand the extent of bias in WTQ, we evaluate LLMs on the no-table baseline. By analyzing the performance of LLMs in the absence of the table, we can better understand the extent to which these models depend on the tabular data.
3 Results
Figure 2: The average EM score for the different few shot comparisons for various perturbations
3.1 Effects of ICL examples on TQA
Does instruction prompting assists LLM for better table comprehension for question answering task?
Fig. 10 compares the performance of various LLMs under different few-shot settings and perturbations, measured by exact match (EM) scores. The first plot illustrates the performances in 0-, 1-, and 2-shot scenarios. Notably, there’s an increasing trend in EM with the number of shots, demonstrating that additional examples improve table comprehension abilities. Thus, we demonstrate complementary results to LLMs on unstructured tasks Wei et al. (2022) in the context of TQA. Fig. 2 focuses on the impact of various perturbations on model performance. Here the models display a drop in EM across different manipulation operations which indicates a sensitivity to changes in the data structure, underscoring a potential vulnerability in model robustness against input invariances and equivariances. The consistency in performance degradation across operations suggests that while models can handle the original task, their structural awareness ability still requires substantial enhancement for broader table comprehension.
3.2 Effects of the Model Type on TQA
Does newer models have better TQA abilities?
Fig 3 shows that Llama3 models generally outperform the Llama2 and Mistral models across different configurations, indicating that newer architectures like Llama3 are more effective at table reasoning tasks. The 70b versions of these models generally perform better than their 7b variants, indicating that larger model sizes enhance reasoning capabilities. Overall, the advancements in model architecture and increased model size significantly contribute to better TQA abilities.
Figure 3: The average EM for different models across the WTQ dataset. The models are categorized by their types (Llama2, Llama3, Mistral) and separated by their sizes (7b, 70b, etc.)
Does instruction tuning help TQA?
The heat maps in Fig. 4 illustrate the performance of various LLMs, demonstrating that models fine-tuned for instructions or conversation exhibit improved performance. For instance, the Llama3-70b-Instruct model significantly outperforms its original version across all table transformations, indicating that instruction-based fine-tuning enhances the model’s ability to handle complex reasoning tasks. Similarly, conversation-focused fine-tuning also leads to better scores, though with a less pronounced improvement compared to instruction-focused tuning. This suggests that fine-tuning models on specific tasks like following instructions or conversing effectively enhances their capability to interpret and manipulate tabular data, making such approaches valuable for improving performance in structured data tasks.
How does model size affect TQA?
Larger models (e.g., Llama3-70b, Mixtral-8x7b) generally show higher performance than smaller models (e.g., Llama2-7b, Mistral-7b), as seen in both the bar plot (Fig 3) and heat maps (Fig 4). For instance, Llama3-70b and Mixtral-8x7b have higher EM scores than Llama2-7b and Mistral-7b. This indicates that model size contributes significantly to TQA performance.
Figure 4: The average EM scores of the original models and fine-tuned for instruction or conversation models across various table augmentation techniques for WTQ dataset
How do performances vary with perturbations?
Table 1 shows the performance of small and large LMs when subjected to various types of value-based perturbation on WTQ. Both model sizes experience a decline in exact match scores across different operations compared to the original setup. The performance with RVP results in a significant performance drop, more so than DVP. This suggests a sharp decrease in the model’s ability to process and comprehend tables when the insertion of arbitrary, non-contextual values compromises the table comprehension ability of LLMs. Conversely, the DVP result indicates that while the model struggles with content that deviates from the original data structure, maintaining data type consistency offers some resilience against total comprehension failure. The observed discrepancies in performance, particularly the deviation from the original model’s performance in scenarios like DVP and RVP, underline a fundamental challenge: these models do not consistently apply their tabular comprehension capabilities when faced with manipulated table values.
VP Operation Large Model Small Model
Original 0.35 0.28
NT 0.06 0.02
DVP 0.20 0.16
RVP 0.13 0.11
NVP 0.06 0.02
Table 1: The average EM of small and large LLMs on Value Perturbations(VP). The classification of large and small models is defined in Table 3
3.3 Effects of Domain Specificity on TQA
How biased are LLMs towards Wiki-tables?
The Table 1 outlines the reliance of small and large language models on structured tabular data, particularly evident in the NT and NVP operations. When no table is provided, both models show a notable performance decline, emphasizing their dependence on tabular data to generate correct answers. Interestingly, the models still manage to answer about $\approx$5% of queries correctly, indicating a potential bias towards Wiki-data. In the NVP scenario, where table values relevant to queries are nullified, there is also a significant drop in performance, yet less severe compared to the complete absence of a table. This suggests that the presence of the original structured format without the relevant data provides enough contextual cues for the models to answer, highlighting model biases.
How do LLMs perform on (TAT-QA), which is a specialized dataset that requires stronger reasoning abilities?
Both large and small LLMs (Fig 5) exhibit moderate performance across various table augmentations on the TAT-QA dataset. Notably, larger models (e.g., Llama3-70b, Mixtral-8x7b) consistently outperform smaller models (e.g., Llama2-7b, Mistral-7b), underscoring their superior TQA abilities. However, the overall scores still denote significant challenges inherent in the dataset. Fine-tuned models, such as Llama3-70b-Instruct, demonstrate enhanced effectiveness, reflecting the advantages of model-specific optimization for harder tabular comprehension tasks. The comparison between the TAT-QA and WTQ datasets highlights a critical insight: while models achieve higher accuracy on WTQ, their moderate EM scores on TAT-QA indicates lack of robustness in TQA. Particularly the sharp contrast in the NT performance, where TAT-QA models have approximately 0% accuracy, a distinguishable contrast from the WTQ performance. This suggests that performance metrics on WTQ might be inflated due to biases favoring familiarity with Wikipedia’s structured data, whereas TAT-QA, a more specialized rarer domain is more challenging. This highlights the need for better benchmark datasets to accurately assess LLMs’ robust table comprehension abilities.
Figure 5: The average EM scores of the original models and fine-tuned for instruction or conversation models across various table augmentation techniques for TAT-QA dataset
4 Conclusion
In conclusion, our study explores the table comprehension abilities of LLMs through the lens of various invariant and equivariant perturbations. We find that instruction tuning and ICL examples significantly enhances table comprehension ability of LLMs. Newer models, such as Llama3, exhibit robust performance compared to their predecessors. Additionally, we uncover domain biases for evaluating tabular data using Wikipedia-based datasets. Despite the decent performance of LLMs on TQA, they still fall short of the reliability required for several practical applications.
5 Limitation
Although we provide extensive evaluation of LLMs on WTQ and TAT-QA datasets, it is possible to including a broader range of datasets for a more comprehensive comparison that would highlight the generalizability of our method for both domain specific dataset and Wikipedia-based dataset. While we anticipate that similar performance could be achieved with other tasks, such as table summarization, future work should include extensive analysis across various tasks and datasets to validate the assumption. Moreover, our study did not involve any structural aware or fine-tuned models for tabular dataset. It is plausible that fine-tuning and structurally enhanced models could significantly impact the performance of different models. Additionally, our evaluation relied on exact match accuracy for assessing the text generation model’s performance. This metric, while useful, limits the scope of evaluation for question answering task. Future studies should employ more nuanced evaluation metrics to better assess the robustness of the models in TQA tasks.
Appendix A Appendix
A.1 Example Prompt
Example of the prompts:
Figure 6: Example of a prompt with answer for WTQ dataset without Few Shot Prompt
Figure 7: Example of a prompt with answer for WTQ dataset for Data Type Preserving Perturbation. In comparison to Fig 6, we replace correct answer(Bankok, Thailand) with a fake answer(Beijing).
Figure 8: Example of a prompt with answer for WTQ dataset for Random Value Perturbation. Here, we replace correct answer(Bankok, Thailand) with an abstract random value (r@nD0m v@1u3).
Figure 9: Example of a prompt with answer for WTQ dataset for Random Value Perturbation. We remove the correct answer(Bankok, Thailand).
Appendix B Evaluation Dataset Size
We select two different datasets for comparison, WTQ and TAT-QA dataset, with different number of evaluation dataset as described in Table 2. For fair comparison we limit the number of cell elements(<150absent150<150< 150) within the table for both datasets. Similarly, for Value Perturbation, some queries relate to over all structure of the table, hence we filter only those tables that contains the answer value for the given query.
Operation Number of Pairs
WTQ dataset
Row Swap 204
Column Swap 204
Transpose 204
Transpose Row Swap 204
Transpose Column Swap 204
Data Type Preserving 141
Random Value 141
Null Value 141
No Table 204
TAT-QA dataset
Row Swap 1668
Column Swap 1668
Transpose 1668
Transpose Row Swap 1668
Transpose Column Swap 1668
No Table 1668
Table 2: Size of the Evaluation datasize
Appendix C Models
We selected recent open source models that have been extensively studied. Table 3 lists all the models we considered with their parameter size and their date of release.
Model Size Date Released
Small Model
Llama-2-7b-hf 6.74B July 2023
Llama-2-7b-chat-hf 6.74B July 2023
Mistral-7B-v0.1 7.24B Sept 2023
Mistral-7B-Instruct-v0.1 7.24B Sept 2023
Meta-Llama-3-8B 8.03B April 2024
Meta-Llama-3-8B-Instruct 8.03B April 2024
Large Model
Llama-2-70b-hf 69B July 2023
Llama-2-70b-chat-hf 69B July 2023
Mixtral-8x7B-v0.1 46.7B Dec 2023
Mixtral-8x7B-Instruct-v0.1 46.7B Dec 2023
Meta-Llama-3-70B 70.6B April 2024
Meta-Llama-3-70B-Instruct 70.6B April 2024
Table 3: All the models with their parameter size and their date released.
Figure 10: The average EM score for the different few shot comparisons for models on the WTQ dataset