### v2

# On the Robustness of Language Models for Tabular Question Answering

Abstract
Large Language Models (LLMs), already shown to ace various text comprehension tasks have also remarkably been shown to tackle table comprehension tasks without specific training. While previous research has explored LLM capabilities with tabular dataset tasks, our study assesses the influence of in-context learning, model scale, instruction tuning, and domain biases on Tabular Question Answering (TQA). We evaluate the robustness of LLMs on Wikipedia-based WTQ, financial report-based TAT-QA, and scientific claims-based SCITAB, TQA datasets, focusing on their ability to interpret tabular data under various augmentations and perturbations robustly. Our findings indicate that instructions significantly enhance performance, with recent models exhibiting greater robustness over earlier versions. However, data contamination and practical reliability issues persist, especially with WTQ. We highlight the need for improved methodologies, including structure-aware self-attention mechanisms and better handling of domain-specific tabular data, to develop more reliable LLMs for table comprehension.
On the Robustness of Language Models for Tabular Question Answering
[…]
1 Introduction
LLMs, primarily trained on unstructured data, have performed fairly well with structured data like tabular datasets. Tasks that rely on tabular structures require the ability to process and understand data with structural awareness, even when the input is provided in a sequential linear format. Exploring such capability—particularly how LLMs apprehend and interpret structured information—can be effectively analyzed by studying robustness measures in table comprehension. This approach offers insights into the models’ ability to navigate and interpret the inherent relationships within structured datasets Borisov et al. (2023).
Figure 1: Example of different possible table augmentation methods where each row operation represents original table augmentation, transpose-based table augmentation, and value-based perturbation. The value-based perturbation only affects the cell content associated with the answer value of the given question.
Tabular Question Answering (TQA) is a popular task on tables, which involves answering a query based on the given structured data Herzig et al. (2020); Fang et al. (2024). TQA task highlights the emergent properties of LLMs, demonstrating their capability to comprehend structured data through in-context learning, instruction adherence, and multi-step reasoning Fang et al. (2024); Wei et al. (2022); Wang et al. (2024); Jiang et al. (2023); Dong et al. (2024); Liu et al. (2024). This capability can be quantitatively evaluated by measuring the robustness of the model’s performance on tabular datasets. Wang et al. (2022) recently showed the vulnerability of LLMs to structural augmentation of the input table and introduced a novel methodology to leverage structure-aware self-attention of an encoder-decoder model like T5 to more effectively manage table transformations, thereby improving the performance of natural language generation from tables. Similarly, Zhao et al. (2023) identified the vulnerabilities of recent TQA LLMs when faced with human-crafted adversarial perturbations. Liu et al. (2023) discusses inherent limitations within LLMs for tabular data interpretation, arguing that these models struggle with structural perturbations and advocate for an integrated approach that combines textual and symbolic reasoning. Chen (2023) find the performance of few-shot reasoning with LLMs for TQA show strong, complex reasoning abilities over table structures without table-specific pre-training.
Although these studies showcase a critical need for enhanced methodologies in handling tabular data across various applications, they have yet to fully explore how the robustness of performance with the inherent ability of LLMs varies with different factors. Notably, there has been little comparison of robustness as a function of (1) in-context learning, (2) the scale and instruction-tuning of LLMs, and (3) the biases inherent within existing evaluation benchmarks for robust table-to-text generation. These aspects are crucial, as the scale of LLMs can influence their ability to generalize from data, while instruction-tuning enhances in-context learning ability, affecting how well they adapt to table comprehension without explicit re-training. Additionally, biases in training data can skew the evaluation of their comprehension ability. Hence, understanding these dimensions is essential for developing more reliable and fair LLMs in processing tabular data Zhou et al. (2024). To evaluate these, Wikipedia-based datasets are the most popular benchmark datasets for TQA. We experiment with WTQ dataset Pasupat and Liang (2015), a Wikipedia-based tabular question-answering dataset, TAT-QA Zhu et al. (2021), a tabular question-answering dataset derived from financial reports, and SCITAB Lu et al. (2023) a scientific tabular question-answering dataset containing challenging claims that require compositional reasoning over tables. A general and specialized domain ensures reliable benchmarking of the robustness and generalizability of existing LLMs.
2 Various Perturbation Categories
Each perturbation is designed to manipulate table structure or content while preserving the inherent relational meaning of the table, to measure robustness to table comprehension, illustrated in Fig. 1.
2.1 Structural Perturbation (SP)
SP involves rearranging the columns and rows of the table to generate new examples. This ensures flexibility in understanding tabular data without distorting the semantics of the table. SP involves column swap, row swap, transpose, transpose column swap, and transpose row swap.
2.2 Value Perturbation (ValP)
ValP focuses on modifying the actual data values within tables, ensuring that the model is faithful to the table. We explore these types of ValP:
Data Type Preserving Perturbation (DVP) [^1] Example of prompts and details of DVP dataset generation are present in the Appendix A, B: DVP involves altering the answer to the question and, respectively, the cell values within the table while maintaining their original data types. For instance, in Fig 6, given a question, ‘‘What was the first venue for the Asian games?’’, we modify the correct answer, ‘‘Bangkok, Thailand", to ‘‘Beijing’’. We use such counterfactual entities to test the faithfulness of the LLM to the tabular data. 222 We filter out the subset of data points where the table does not contain the answer, e.g., ‘‘How many people stayed at least 3 years in office?’’. Random Value Perturbation (RVP): RVP (an example is shown in Fig. 7) relaxes DVP where instead of a counterfactual entity, we have a fixed string, e.g., ‘‘r@nD0m v@1u3’’.
Null Value Perturbation (NVP): NVP removes the correct answer from the table completely. Evaluating the performance on NVP highlights the influence of Wikipedia content on solving the WTQ table question-answering task.
No Table (NT): To understand the extent of bias in WTQ, we evaluate LLMs on the no-table baseline. By analyzing the performance of LLMs in the absence of the table, we can better understand the extent of dependence on the particular tabular data.
3 Results
Figure 2: The average EM score across three different datasets for the different few shot comparisons for various perturbations with three different datasets.
3.1 Effects of ICL examples on TQA
Does instruction prompting assist LLM for better table comprehension for question-answering tasks? Fig. 11 compares the performance of various LLMs under different few-shot settings, measured by exact match (EM) scores. The plot illustrates the performances in 0, 1, 2, and 3-shot scenarios. Notably, there’s an increasing trend in EM with the number of shots, demonstrating that additional examples improve table comprehension abilities. Following the patterns observed in the WTQ dataset, the overall performance trends for both TAT-QA (Fig.13) and SCITAB (Fig.14) are generally consistent. However, within SCITAB, there are a few exceptions where smaller models outperform their larger counterparts. In general, we demonstrate complementary results to LLMs on unstructured tasks Wei et al. (2022) in the context of TQA. Fig. 2 and 12 focuses on the impact of various perturbations on model performance. Here, the models display a drop in EM across different manipulation operations, which indicates a sensitivity to changes in the data structure, underscoring a potential vulnerability in model robustness against input invariances and equivariances. The consistency in performance degradation across operations suggests that while models can handle the original task, their structural awareness ability still requires substantial enhancement for broader table comprehension.
3.2 Effects of the Model Type on TQA
Does newer models have better TQA abilities? Fig. 3 shows that Llama3 models generally outperform the Llama2 and Mistral models across different configurations, indicating that newer architectures like Llama3 are more effective at table reasoning tasks. The 70b versions of these models generally perform better than their 7b variants, indicating that larger model sizes enhance reasoning capabilities. Overall, the advancements in model architecture and increased model size significantly contribute to better TQA abilities.
Figure 3: The average EM for different models across the WTQ dataset with 3-fewshot settings. The models are categorized by their types (Llama2, Llama3, Mistral) and separated by their sizes (7b, 70b, etc.)
Does instruction tuning help TQA? The heat maps in Fig. 4 illustrate the performance of various LLMs, demonstrating that models fine-tuned for instructions or conversation exhibit improved performance. For instance, the Llama3-70b-Instruct model significantly outperforms its original version across all table transformations, indicating that instruction-based fine-tuning enhances the model’s ability to handle complex reasoning tasks. Similarly, conversation-focused fine-tuning also leads to better scores, though with a less pronounced improvement compared to instruction-focused tuning. This suggests that fine-tuning models on specific tasks like following instructions or conversing effectively enhances their capability to interpret and manipulate tabular data, making such approaches valuable for improving performance in structured data tasks. Figure 4(right) distinctly indicate that LLMs that have undergone instruction or conversation-based fine-tuning outperform their base counterparts in TQA tasks for SCITAB dataset. Although SCITAB is inherently a classification task presented in a TQA format, using few-shot prompting provides valuable context, ultimately leading to more accurate and relevant responses.
How does model size affect TQA? Larger models (e.g., Llama3-70b, Mixtral-8x7b) generally show higher performance than smaller models (e.g., Llama2-7b, Mistral-7b), as seen in both the bar plot (Fig 3) and heat maps (Fig 4). For instance, Llama3-70b and Mixtral-8x7b have higher EM scores than Llama2-7b and Mistral-7b. This indicates that model size contributes significantly to TQA performance.
Figure 4: The average EM scores of the original models and fine-tuned for instruction or conversation models across various table augmentation techniques for WTQ, TAT-QA and SCITAB dataset
How do performances vary with value perturbations?  [^3] More details are in Appendix Section F and G Table 1 and 4 show the performance of small and large LLMs, as defined on Table 3, when subjected to various types of value-based perturbation on WTQ using different evaluation metrics, Exact Match(EM), Variation percentage (VP) Yang et al. (2022); Zhou et al. (2024), and Exact Match Difference (Emd)Zhao et al. (2023); Zhou et al. (2024). Both model sizes experience a decline in exact match scores across different operations compared to the original setup. The performance with RVP results in a significant performance drop, more so than DVP. This suggests a sharp decrease in the model’s ability to process and comprehend tables when the insertion of arbitrary, non-contextual values compromises the table comprehension ability of LLMs. Conversely, the DVP result indicates that while the model struggles with content that deviates from the original data structure, maintaining data type consistency offers resilience against total comprehension failure. VP increases significantly across all perturbations for both large and small models, indicating considerable changes in predictions due to the perturbations. EMD consistently shows negative values, with the most substantial performance drops occurring in the NT and NVP scenarios, reflecting the models’ vulnerability to these disruptions. The observed discrepancies in performance, particularly the deviation from the original model’s performance in scenarios like DVP and RVP, underline a fundamental challenge: these models do not consistently apply their tabular comprehension capabilities when faced with manipulated table values.
ValP Large Model Small Model
Operation EM VP Emd EM VP Emd
Original 0.36 - - 0.25 - -
NT 0.05 0.36 -0.31 0.02 0.25 -0.23
DVP 0.21 0.40 -0.14 0.14 0.31 -0.11
RVP 0.14 0.39 -0.20 0.10 0.30 -0.15
NVP 0.07 0.35 -0.28 0.03 0.26 -0.23
Table 1: The average EM, VP, and Emd of small and large LLMs on Value Perturbations(ValP). The classification of large and small models is defined in Table 3.
3.3 Effects of Domain Specificity on TQA
How biased are LLMs towards Wiki-tables? Table 1 outlines the reliance of LLMs on structured tabular data, particularly evident in the NT and NVP operations. When no table is provided, both type of models show a notable performance decline, emphasizing their dependence on tabular data to generate correct answers. Interestingly, the models still manage to answer about $\approx$5of queries correctly, indicating a potential bias towards Wiki-data. In the NVP scenario, where table values relevant to queries are nullified, there is a significant drop in performance, yet less severe compared to the complete absence of a table. This suggests that the original structured format without the relevant data provides enough contextual cues for the models to answer, highlighting model biases.
How do LLMs perform on (TAT-QA) and (SCITAB), which is a specialized dataset that requires stronger reasoning abilities? Both large and small LLMs exhibit moderate performance across various table augmentations on the TAT-QA dataset(Fig 4(middle)) and SCITAB dataset(Fig 4(right)). For TAT-QA, large models consistently outperform smaller models, underscoring their superior TQA abilities, but for SCITAB, the small models have better performance in comparison to larger models. The overall inconsistent scores still denote significant challenges inherent in the domain-specific dataset. Instruction-tuned models, demonstrate enhanced effectiveness, reflecting the advantages of model-specific optimization for harder tabular comprehension tasks. From Table 4, the comparison among WTQ, TAT-QA, and SCITAB datasets reveals a critical insight: while models start with higher EM on WTQ but suffer larger Emd and higher VP, indicating higher sensitivity to perturbations, their relatively smaller Emd and lower VP on both TAT-QA and SCITAB reflect greater robustness—despite SCITAB ’s overall lower baseline EM.. Particularly the sharp contrast in the NT performance, where on TAT-QA and SCITAB dataset, models have approximately 0% accuracy, a distinguishable contrast from the WTQ performance. This suggests that performance metrics on WTQ might be inflated due to biases favoring familiarity with Wikipedia’s structured data, whereas TAT-QA and SCITAB, a more specialized rarer domain, is more challenging. This highlights the need for better benchmark datasets to assess LLMs’ robust table comprehension abilities accurately.
4 Conclusion
Our study explores the table comprehension abilities of LLMs through the lens of various invariant and equivariant perturbations. We find that instruction tuning and ICL examples enhance the table comprehension ability of LLMs. Newer models, such as Llama3, are more robust compared to their predecessors. Additionally, we uncover domain biases for evaluating tabular data using Wikipedia-based datasets. Despite the decent performance of LLMs on TQA, they still fall short of the reliability required for several practical applications.
5 Limitation
Although we provide extensive evaluation of LLMs on WTQ, TAT-QA, and SCITAB datasets, it is possible to including a broader range of datasets for a more comprehensive comparison that would highlight the generalizability of our method for both domain specific dataset and Wikipedia-based dataset. While we anticipate that similar performance could be achieved with other tasks, such as table summarization, future work should include extensive analysis across various tasks and datasets to validate the assumption. Moreover, our study did not involve any structural aware or fine-tuned models for tabular dataset. It is plausible that fine-tuning and structurally enhanced models could significantly impact the performance of different models. Additionally, our evaluation relied on exact match accuracy for assessing the text generation model’s performance. This metric, while useful, limits the scope of evaluation for question answering task. Future studies should employ more nuanced evaluation metrics to better assess the robustness of the models in TQA tasks.
Appendix A Example Prompt
Example of the prompts:
Figure 5: Example of a prompt with answer for WTQ dataset without Few Shot Prompt
Figure 6: Example of a prompt with answer for WTQ dataset for Data Type Preserving Perturbation. In comparison to Fig 5, we replace correct answer(Bankok, Thailand) with a fake answer(Beijing).
Figure 7: Example of a prompt with answer for WTQ dataset for Random Value Perturbation. Here, we replace correct answer(Bankok, Thailand) with an abstract random value (r@nD0m v@1u3).
Figure 8: Example of a prompt with answer for WTQ dataset for Random Value Perturbation. We remove the correct answer(Bankok, Thailand).
Appendix B Generating Data Type Preserving Perturbation Dataset
We utilize ChatGPT-3.5 to generate type-preserving counterfactual answers using the prompt illustrated in Fig. 9. The prompt is designed to ensure that any generated answer aligns with the data type of the original answer while deliberately introducing a perturbation. For instance, given a question like "What was the first venue for the Asian games?" with the correct answer "Bangkok, Thailand," the model is prompted to output a different but type-consistent response, such as "Beijing." Similarly, for numerical data, a value like "42" might be replaced with another plausible number, such as "50," ensuring that the altered answer retains the original datatype constraints. This is achieved by explicitly defining the model’s role as a generator of "fake" answers that preserve the data type of the original value. To ensure the validity of value perturbations, we select only tables that contain the exact answer to the given question. These constraints guarantee that data type preserving perturbations are applied exclusively to table-question-answer triples where the table explicitly contains the correct answer required to solve the question.
Figure 9: Prompt designed for generating one-word, datatype-matching fake answers to questions from the WTQ dataset.
Appendix C Evaluation Dataset Size
We select three different datasets for comparison, WTQ, TAT-QA, and SCITAB dataset, with different numbers of evaluation datasets as described in Table 2. For fair comparison, we limit the number of cell elements($<150$) within the table for both datasets. Similarly, for Value Perturbation, some queries relate to over all structure of the table, hence we filter only those tables that contains the answer value for the given query.
Operation Number of Pairs
WTQ dataset
Row Swap 204
Column Swap 204
Transpose 204
Transpose Row Swap 204
Transpose Column Swap 204
Data Type Preserving 141
Random Value 141
Null Value 141
No Table 204
TAT-QA dataset
Row Swap 1668
Column Swap 1668
Transpose 1668
Transpose Row Swap 1668
Transpose Column Swap 1668
No Table 1668
SCITAB dataset
Row Swap 1225
Column Swap 1225
Transpose 1225
Transpose Row Swap 1225
Transpose Column Swap 1225
No Table 1225
Table 2: Size of the Evaluation datasize
Appendix D Models
We selected recent open-source models that have been extensively studied. Table 3 lists all the models we considered with their parameter size and their date of release.
Model Size Date Released
Small Model(Less than 10B parameter)
Llama-2-7b-hf 6.74B July 2023
Llama-2-7b-chat-hf 6.74B July 2023
Mistral-7B-v0.1 7.24B Sept 2023
Mistral-7B-Instruct-v0.1 7.24B Sept 2023
Meta-Llama-3-8B 8.03B April 2024
Meta-Llama-3-8B-Instruct 8.03B April 2024
Large Model(Larger than 40B parameter)
Llama-2-70b-hf 69B July 2023
Llama-2-70b-chat-hf 69B July 2023
Mixtral-8x7B-v0.1 46.7B Dec 2023
Mixtral-8x7B-Instruct-v0.1 46.7B Dec 2023
Meta-Llama-3-70B 70.6B April 2024
Meta-Llama-3-70B-Instruct 70.6B April 2024
Table 3: All the models with their parameter size and their date released. Large Models are defined as models with parameters larger than 40 billion parameters, and Small Models are models with parameters smaller than 10 billion parameters.
Appendix E Evaluation Metrics
We employ three primary metrics to evaluate model performance under various perturbations in Tabular Question Answering (TQA) tasks.
Exact Match Accuracy (EM): This metric calculates the proportion of instances for which the predicted answer exactly matches the ground truth answer. Formally, if $N$ is the total number of instances, and $\text{correct}(i)$ is an indicator function that is $1$ if the prediction for instance $i$ is correct and 00 otherwise, then:
$\text{EM}=\frac{\sum_{i=1}^{N}\text{correct}(i)}{N}.$
Exact Match Difference (Emd): Let $\text{EM}_{\text{orig}}$ be the EM on the original (unperturbed) dataset, and $\text{EM}_{\text{perturbed}}$ be the EM after applying a perturbation. The Emd quantifies the change in EM due to perturbations:
$\text{Emd}=\text{EM}_{\text{perturbed}}-\text{EM}_{\text{orig}}.$
Negative values indicate performance degradation, while values close to zero imply robustness against perturbations.
Variation Percentage (VP): This metric measures to what extent predictions change after applying perturbations. Given that, C2W is the count of correct before perturbation and wrong after, and W2C is the count of wrong before perturbation and correct after. Given $N$ as the total number of instances, the variation percentage is:
$\text{VP}=\frac{C2W+W2C}{N}.$
A higher VP indicates greater sensitivity of predictions to perturbations, while a lower VP signifies more stable predictions.
Appendix F Performance over Table Complexity
Figure 10 presents average model performance across varying table sizes for three tabular question-answering datasets—WTQ, TAT-QA, and SCITAB—under several table structure perturbations. Each plot compares the original condition with different table augmentation operations, such as column swapping, row swapping, and transpose-based modifications.
For WTQ, performance is low overall and highly variable across increasing table sizes, with no single condition consistently outperforming others. In contrast, TAT-QA shows a generally higher and more stable baseline, though still affected by growing table size and perturbations; performance tends to decline as tables grow larger, suggesting sensitivity to complexity. SCITAB results are somewhat mixed, with fluctuations at different size intervals, but the performance remains closer among the different conditions.
Across all three datasets, these results highlight that large language models, though capable, display uneven robustness to structural manipulations of tables. Although instruction tuning and larger model scales improve performance, structural changes continue to pose challenges, indicating the need for more robust, structure-aware approaches for reliable table comprehension.
Figure 10: The average EM score for the different structural perturbation over the different size of tables on WTQ, TAT-QA and SCITAB datasets
Appendix G Additional Robustness Result
Figure 11: Comparison of average EM scores across models under varying few-shot settings (0-shot, 1-shot, 2-shot, and 3-shot).
Figure 12: The average cumulative EM score across three different datasets for the different few shot comparisons for various perturbations with three different datasets.
The table 4 summarizes the impact of various perturbation operations on the accuracy and robustness of larger and smaller LLMs models when performing TQA across three distinct datasets (WTQ, TAT-QA, and SCITAB). For each perturbation categories the EM scores typically drop compared to the original configuration, as indicated by negative EmD values. This performance dip is especially notable on the WTQ dataset, where even minor perturbations lead to reduced EM, and certain operations like column transpose and NT conditions cause pronounced declines in EM. Meanwhile, the VP metrics show that predictions shift more substantially under these transformations, signaling sensitivity in both large and small models to altered table structures. Although large models usually have higher baseline EM, they still exhibit performance degradation under perturbations. Particularly striking are conditions like NT—where the table is effectively removed—resulting in EM plummeting to near-zero and high VP, underscoring the models’ heavy reliance on the given tabular data. The results highlight the need for improved architectural and training strategies aimed at enhancing structural awareness and mitigating the negative effects of domain-specific perturbations, thereby making LLMs more robust and reliable in real-world table comprehension scenarios.
Operation WTQ Dataset TATQA Dataset SCITAB Dataset
Large Model Small Model Large Model Small Model Large Model Small Model
EM VP Emd EM VP Emd EM VP Emd EM VP Emd EM VP Emd EM VP Emd
Original 0.36 - - 0.25 - - 0.33 - - 0.20 - - 0.098 - - 0.132 - -
Column 0.35 0.16 -0.01 0.25 0.15 0.00 0.31 0.09 -0.03 0.17 0.09 -0.03 0.073 0.085 -0.026 0.131 0.080 -0.001
Row 0.30 0.19 -0.06 0.20 0.17 -0.05 0.30 0.11 -0.03 0.18 0.09 -0.03 0.082 0.080 -0.017 0.132 0.079 0.000
Transpose 0.32 0.19 -0.04 0.20 0.18 -0.05 0.31 0.09 -0.02 0.18 0.09 -0.02 0.085 0.086 -0.015 0.134 0.082 0.002
Col Transpose 0.24 0.23 -0.12 0.15 0.20 -0.10 0.29 0.11 -0.05 0.16 0.10 -0.04 0.085 0.084 -0.013 0.130 0.082 -0.001
Row Transpose 0.31 0.20 -0.05 0.20 0.18 -0.05 0.26 0.13 -0.07 0.14 0.11 -0.06 0.088 0.085 -0.009 0.132 0.082 0.001
NT 0.06 0.36 -0.31 0.02 0.25 -0.23 0.01 0.33 -0.32 0.00 0.20 -0.20 0.000 0.113 -0.098 0.000 0.132 -0.132
DVP 0.21 0.40 -0.14 0.14 0.31 -0.11 - - - - - - - - - - - -
RVP 0.14 0.39 -0.20 0.10 0.30 -0.15 - - - - - - - - - - - -
NVP 0.07 0.35 -0.28 0.03 0.26 -0.23 - - - - - - - - - - - -
Table 4: The average EM, VP, and Emd of small and large LLMs on all perturbations operations
Figure 13: The average EM for different models across the TAT-QA dataset with 3-fewshot settings. The models are categorized by their types (Llama2, Llama3, Mistral) and separated by their sizes (7b, 70b, etc.)
Figure 14: Average EM scores for both original and instruction or conversation fine-tuned models on the SCITAB dataset (3-shot setting) under various table augmentation techniques.