### v3

# Exploring the Robustness of Language Models for Tabular Question Answering via Attention Analysis

Abstract
Large Language Models (LLMs), already shown to ace various text comprehension tasks, have also remarkably been shown to tackle table comprehension tasks without specific training. Building on earlier studies of LLMs for tabular tasks, we probe how in-context learning (ICL), model scale, instruction tuning, and domain bias affect Tabular QA (TQA) robustness by testing LLMs, under diverse augmentations and perturbations, on diverse domains: Wikipedia-based WTQ, financial TAT-QA, and scientific SCITAB. Although instruction tuning and larger, newer LLMs deliver stronger, more robust TQA performance, data contamination and reliability issues, especially on WTQ, remain unresolved. Through an in-depth attention analysis, we reveal a strong correlation between perturbation-induced shifts in attention dispersion and the drops in performance, with sensitivity peaking in the model’s middle layers. We highlight the need for improved interpretable methodologies to develop more reliable LLMs for table comprehension.
Exploring the Robustness of Language Models for Tabular Question Answering via Attention Analysis
[…]
footnotetext: * These authors jointly supervised this work.
1 Introduction
LLMs, despite being primarily trained on unstructured text, have demonstrated notable capabilities in structured-data tasks such as Tabular Question Answering (TQA). TQA requires models to interpret data presented in tables, demanding strong structural reasoning. Assessing how LLMs navigate structured comprehension challenges can provide valuable insights into their robustness and reasoning capabilities Borisov et al. (2023); Fang et al. (2024).
Recent studies highlight the importance of robustness evaluations in understanding LLM behavior on structured tasks Zhou et al. (2024). Specifically, perturbations in tabular structure and content significantly impact model performance, highlighting vulnerabilities that limit practical reliability Wang et al. (2022); Zhao et al. (2023). Furthermore, Liu et al. (2023) argue that LLMs inherently struggle with structural manipulations, advocating for an integrated approach combining symbolic reasoning to enhance model robustness.
Figure 1: (a) An example of the different possible table augmentation methods.(b) Scatter plot between change in attention entropy and EM score across perturbation types for Llama-8B-Instruct on WTQ.
Although these studies identify vulnerabilities and suggest potential improvements, there remains limited understanding of how internal model mechanisms respond to perturbations in structured data. Attention mechanisms form the core of transformer-based LLMs, governing how models distribute focus across input elements during processing Clark et al. (2019). Prior work analyzing attention patterns in natural language tasks revealed that specific layers and attention heads critically influence model performance and robustness Zhao et al. (2024); Barbero et al. (2025). However, detailed attention-level analyses for structured tasks like TQA remain scarce.
To address this gap, we systematically investigate LLM robustness for TQA tasks across multiple dimensions, including in-context learning, model scale, instruction tuning, domain biases, value perturbations, and model size. Our experiments evaluate different perturbation types, highlighting how value-based alterations influence reasoning fidelity and faithfulness (Table 1). We also compare performance across diverse datasets to assess domain biases and generalizability.
Extending beyond surface-level performance metrics, we quantify changes in attention entropy across different attention heads and layers under various perturbations, analyzing how these changes correlate with performance degradation. With this we aim to uncover which attention heads are most sensitive to perturbations, informing targeted improvements for future models.
Our study leverages diverse datasets: Wikipedia-based WTQ Pasupat and Liang (2015), financial report-based TAT-QA Zhu et al. (2021), and scientific claims-based SCITAB Lu et al. (2023). Through these diverse domains, our findings give insights into the sensitivity and reliability of LLMs on TQA, highlighting the broader challenge of understanding how LLMs reason over structured data.
2 Various Perturbation Categories
Each perturbation is designed to manipulate table structure or content while preserving the inherent relational meaning of the table, to measure robustness to table comprehension, illustrated in Fig. 1. Structural Perturbation (SP): SP involves rearranging the columns and rows of the table to generate new examples. This ensures flexibility in understanding tabular data without distorting the semantics of the table. SP involves column swap, row swap, transpose, transpose column swap, and transpose row swap.
Value Perturbation (ValP): ValP focuses on modifying the actual data values within tables, ensuring that the model is faithful to the table. We explore these types of ValP:
Data Type Preserving Perturbation (DVP) [^1] Example of prompts and details of DVP dataset generation are present in the Appendix A, B: DVP involves altering the answer to the question and, respectively, the cell values within the table while maintaining their original data types. For instance, in Fig 5, given a question, ‘‘What was the first venue for the Asian games?’’, we modify the correct answer, ‘‘Bangkok, Thailand", to ‘‘Beijing’’. These counterfactual entities test the faithfulness of the LLM to the tabular data. 222 We filter out the subset of data points where the table does not contain the answer, e.g., ‘‘How many people stayed at least 3 years in office?’’.Random Value Perturbation (RVP): RVP (an example is shown in Fig. 6) relaxes DVP where instead of a counterfactual entity, we have a fixed string, e.g., ‘‘r@nD0m v@1u3’’.
Null Value Perturbation (NVP): NVP removes the correct answer from the table completely. Evaluating the performance on NVP highlights the influence of Wikipedia content on solving the WTQ table question-answering task.
No Table (NT): To understand the extent of bias in WTQ, we evaluate LLMs on the no-table baseline. By analyzing the performance of LLMs in the absence of the table, we can better understand the extent of dependence on the particular tabular data.
3 Evaluation Performance
Operation WTQ Dataset TATQA Dataset SCITAB Dataset
Large Model Small Model Large Model Small Model Large Model Small Model
EM VP Emd EM VP Emd EM VP Emd EM VP Emd EM VP Emd EM VP Emd
Original 0.36 - - 0.25 - - 0.33 - - 0.20 - - 0.098 - - 0.132 - -
Column 0.35 0.16 -0.01 0.25 0.15 0.00 0.31 0.09 -0.03 0.17 0.09 -0.03 0.073 0.085 -0.026 0.131 0.080 -0.001
Row 0.30 0.19 -0.06 0.20 0.17 -0.05 0.30 0.11 -0.03 0.18 0.09 -0.03 0.082 0.080 -0.017 0.132 0.079 0.000
Transpose 0.32 0.19 -0.04 0.20 0.18 -0.05 0.31 0.09 -0.02 0.18 0.09 -0.02 0.085 0.086 -0.015 0.134 0.082 0.002
Col Transpose 0.24 0.23 -0.12 0.15 0.20 -0.10 0.29 0.11 -0.05 0.16 0.10 -0.04 0.085 0.084 -0.013 0.130 0.082 -0.001
Row Transpose 0.31 0.20 -0.05 0.20 0.18 -0.05 0.26 0.13 -0.07 0.14 0.11 -0.06 0.088 0.085 -0.009 0.132 0.082 0.001
NT 0.06 0.36 -0.31 0.02 0.25 -0.23 0.01 0.33 -0.32 0.00 0.20 -0.20 0.000 0.113 -0.098 0.000 0.132 -0.132
DVP 0.21 0.40 -0.14 0.14 0.31 -0.11 - - - - - - - - - - - -
RVP 0.14 0.39 -0.20 0.10 0.30 -0.15 - - - - - - - - - - - -
NVP 0.07 0.35 -0.28 0.03 0.26 -0.23 - - - - - - - - - - - -
Table 1: The average EM, VP, and Emd of small and large LLMs on all perturbation operations. Large and Small LLMs are defined on Table 3.
3.1 Effects of the Model Type on TQA
How do performances vary with value perturbations?  [^3] More details are in Appendices F and G Table 1 shows the performance of various LLMs, as defined on Table 3, when subjected to different value-based perturbation on WTQ using different evaluation metrics: Exact Match(EM), Variation percentage (VP) Yang et al. (2022); Zhou et al. (2024), and Exact Match Difference (EMD)Zhao et al. (2023); Zhou et al. (2024). We see LLMs experience a decline in EM scores across different operations compared to the original setup. The performance with RVP results in a significant performance drop, more so than DVP. This suggests a sharp decrease in the model’s ability to process and comprehend tables when the insertion of arbitrary, non-contextual values compromises the table comprehension ability of LLMs. Conversely, the DVP result indicates that while the model struggles with content that deviates from the original data structure, maintaining data type consistency helps. VP increases significantly across all perturbations indicating considerable changes in predictions due to the perturbations. EMD consistently shows negative values, with the most substantial performance drops occurring in the NT and NVP scenarios. The observed discrepancies in performance, particularly pronounced in DVP and RVP, underline a fundamental challenge: these models do not consistently apply their tabular comprehension capabilities when faced with perturbed tables.
3.2 Effects of Domain Specificity on TQA
How biased are LLMs towards Wiki-tables? Table 1 shows that when no table is provided, LLMs show a notable performance decline, emphasizing their dependence on tabular data to generate correct answers. Interestingly, the models still manage to answer about $\approx$5of queries correctly, indicating a potential bias in Wiki-data. In the NVP scenario, where table values relevant to queries are nullified, there is a significant drop in performance, yet less severe compared to the complete absence of a table. This suggests that the models are biased by the contextual structured format cue even in the absence of relevant data.
How do out-of-box LLMs perform on specialized domains: TAT-QA and SCITAB ? As shown in Table 1, LLMs exhibit moderate performance across various table augmentations on the TAT-QA dataset and SCITAB dataset. For TAT-QA, large models consistently outperform smaller models, underscoring their superior TQA abilities, but for SCITAB, the small models have better performance in comparison to larger models. The overall inconsistent scores still denote significant challenges inherent in niche domains, but instruction-tuned models are better here. From Table 1, the comparison among WTQ, TAT-QA, and SCITAB datasets reveals an interesting accuracy-robustness tradeoff: while models with higher EM on WTQ suffer larger EMD and higher VP, indicating higher sensitivity to perturbations, their relatively smaller EMD and lower VP on both TAT-QA and SCITAB reflect greater robustness, despite SCITAB ’s overall lower baseline EM. Also notable is the sharp contrast in the NT performance, where on TAT-QA and SCITAB dataset, models have approximately 0% accuracy, a contrast from the WTQ performance. This suggests that performance on WTQ might be inflated due to biases favoring familiarity with Wikipedia, compared to niche domains like TAT-QA and SCITAB. This highlights the need for better benchmark design for tabular understanding. We also observe trends consistent with prior work: performance improves with few-shot prompting, instruction-tuned models generally outperform their base counterparts, and larger models tend to exhibit stronger TQA Wei et al. (2022); Fang et al. (2024). Additional results and full breakdowns across datasets, model sizes, and fine-tuning configurations are included in Appendix G.
4 Attention Analysis
Figure 2: Heatmap showing Spearman correlation between changes in attention entropy and EM difference across all attention heads and layers in the LLama-8B-Instruct model on the WTQ dataset.
Understanding how structural perturbations in tabular inputs affect a model is critical for assessing robustness and diagnosing potential failure modes in TQA. As attention mechanisms regulate how language models allocate focus across table elements, examining their sensitivity to perturbations provides insights into representational stability. Here, we analyze this sensitivity by quantifying how perturbation-induced changes in attention dispersion correlate with performance degradation.
4.1 Effect of Perturbations on Attention Maps
We examine how varying severities of structural perturbations influence attention weights using the Llama3-8B-Instruct model on the WTQ dataset. For each attention head across all layers, we measure the change in attention entropy (detailed in Appendix H) between the original and perturbed table inputs. In parallel, we compute the corresponding change in Exact Match (EM) scores, capturing the performance impact of each perturbation.
Attention entropy captures the distribution uniformity of attention across tokens; higher entropy indicates diffuse attention, while lower entropy signifies focused attention. Thus, significant entropy changes reflect considerable shifts in the model’s internal attentional focus due to perturbations.
Our analysis (Figure 1) reveals a significant positive correlation (Spearman $\rho=0.90$, $p=0.037$) between changes in attention entropy and EM degradation across perturbation types. This shows that more severe perturbations significantly disrupt the model’s attention distribution, consequently diminishing its ability for TQA.
These findings highlight the sensitivity of attention mechanisms to structural integrity within tabular inputs. Perturbations affecting relational semantics induce greater attention dispersion and misalignment, directly impairing model accuracy. Thus, attention dispersion is a crucial indicator linking input perturbations with performance.
4.2 Layer-Wise Analysis of Attention Head Sensitivity to Perturbations
Figure 3: Layer-wise Spearman correlation between aggregated attention entropy change (averaged over heads) and EM difference.
While averaging attention metrics across heads provides a general understanding of model behavior, prior research emphasizes that individual attention heads and specific layers, particularly the middle layers, significantly contribute to encoding task-relevant information Barbero et al. (2025); Zhao et al. (2024). To investigate this at a finer granularity, we perform a layer-wise analysis of how perturbation-induced changes in attention entropy correlate with EM score degradation.
Specifically, we use the Llama3-8B-Instruct model on the WTQ dataset to compute the Spearman correlation between entropy changes (original vs. perturbed inputs) and corresponding EM differences for each attention head across all layers. Figure 2 demonstrates that correlations peak predominantly in the middle layers and remain consistently elevated through these central layers. This indicates that perturbation-induced shifts in attention distribution within middle layers are strongly predictive of performance degradation. Additional experiments are included within the Appendix H.
To further substantiate this pattern, we aggregate entropy changes across all heads within each layer and analyze the correlation with EM differences (Figure 3). This aggregated analysis reinforces the observation of heightened sensitivity in mid-layer attention mechanisms.Collectively, these results highlight that perturbation sensitivity is not uniformly distributed across model layers. Instead, the middle layers, which typically engage in critical representational refinement, are particularly vulnerable to attention instability and associated performance degradation.
5 Conclusion
Our study offers a comprehensive robustness analysis of LLMs under various perturbations for TQA. While larger, instruction-tuned models show improved performance, they remain highly sensitive to structural and value-based disruption. We also uncover domain biases, where models perform well on WTQ without tables yet struggle on more specialized datasets like TAT-QA and SCITAB. We show that shifts in attention entropy, particularly in middle layers, correlate with performance degradation. This work highlights the importance of interpretability and robustness of LLMs for TQA.
6 Limitation
Although we provide extensive evaluation of LLMs on WTQ, TAT-QA, and SCITAB datasets, it is possible to include a broader range of datasets for a more comprehensive comparison that would highlight the generalizability of our method for both domain-specific datasets and Wikipedia-based datasets. While we anticipate that similar performance could be achieved with other tasks, such as table summarization, future work should include extensive analysis across various tasks and datasets to validate the assumption. Moreover, our study did not involve any structurally aware or fine-tuned models for tabular datasets. It is plausible that fine-tuning and structurally enhanced models could significantly impact the performance of different models. Additionally, our evaluation relied on exact match accuracy for assessing the text generation model’s performance. This metric, while useful, limits the scope of evaluation for the question answering task. Future studies should employ more nuanced evaluation metrics to better assess the robustness of the models in TQA tasks. Moreover, we only conduct a case study of attention analysis with small models because of computation cost.
Appendix A Example Prompt
Example of the prompts:
Figure 4: Example of a prompt with answer for WTQ dataset without Few Shot Prompt
Figure 5: Example of a prompt with answer for WTQ dataset for Data Type Preserving Perturbation. In comparison to Fig 4, we replace correct answer(Bankok, Thailand) with a fake answer(Beijing).
Figure 6: Example of a prompt with answer for WTQ dataset for Random Value Perturbation. Here, we replace correct answer(Bankok, Thailand) with an abstract random value (r@nD0m v@1u3).
Figure 7: Example of a prompt with answer for WTQ dataset for Random Value Perturbation. We remove the correct answer(Bankok, Thailand).
Appendix B Generating Data Type Preserving Perturbation Dataset
We utilize ChatGPT-3.5 to generate type-preserving counterfactual answers using the prompt illustrated in Fig. 8. The prompt is designed to ensure that any generated answer aligns with the data type of the original answer while deliberately introducing a perturbation. For instance, given a question like "What was the first venue for the Asian games?" with the correct answer "Bangkok, Thailand," the model is prompted to output a different but type-consistent response, such as "Beijing." Similarly, for numerical data, a value like "42" might be replaced with another plausible number, such as "50," ensuring that the altered answer retains the original datatype constraints. This is achieved by explicitly defining the model’s role as a generator of "fake" answers that preserve the data type of the original value. To ensure the validity of value perturbations, we select only tables that contain the exact answer to the given question. These constraints guarantee that data type preserving perturbations are applied exclusively to table-question-answer triples where the table explicitly contains the correct answer required to solve the question.
Figure 8: Prompt designed for generating one-word, datatype-matching fake answers to questions from the WTQ dataset.
Appendix C Evaluation Dataset Size
We select three different datasets for comparison, WTQ, TAT-QA, and SCITAB dataset, with different numbers of evaluation datasets as described in Table 2. For fair comparison, we limit the number of cell elements($<150$) within the table for both datasets. Similarly, for Value Perturbation, some queries relate to over all structure of the table, hence we filter only those tables that contains the answer value for the given query.
Operation Number of Pairs
WTQ dataset
Row Swap 204
Column Swap 204
Transpose 204
Transpose Row Swap 204
Transpose Column Swap 204
Data Type Preserving 141
Random Value 141
Null Value 141
No Table 204
TAT-QA dataset
Row Swap 1668
Column Swap 1668
Transpose 1668
Transpose Row Swap 1668
Transpose Column Swap 1668
No Table 1668
SCITAB dataset
Row Swap 1225
Column Swap 1225
Transpose 1225
Transpose Row Swap 1225
Transpose Column Swap 1225
No Table 1225
Table 2: Size of the Evaluation datasize
Appendix D Models
We selected recent open-source models that have been extensively studied. Table 3 lists all the models we considered with their parameter size and their date of release.
Model Size Date Released
Small Model(Less than 10B parameter)
Llama-2-7b-hf 6.74B July 2023
Llama-2-7b-chat-hf 6.74B July 2023
Mistral-7B-v0.1 7.24B Sept 2023
Mistral-7B-Instruct-v0.1 7.24B Sept 2023
Meta-Llama-3-8B 8.03B April 2024
Meta-Llama-3-8B-Instruct 8.03B April 2024
Large Model(Larger than 40B parameter)
Llama-2-70b-hf 69B July 2023
Llama-2-70b-chat-hf 69B July 2023
Mixtral-8x7B-v0.1 46.7B Dec 2023
Mixtral-8x7B-Instruct-v0.1 46.7B Dec 2023
Meta-Llama-3-70B 70.6B April 2024
Meta-Llama-3-70B-Instruct 70.6B April 2024
Table 3: All the models with their parameter size and their date released. Large Models are defined as models with parameters larger than 40 billion parameters, and Small Models are models with parameters smaller than 10 billion parameters.
Appendix E Evaluation Metrics
We employ three primary metrics to evaluate model performance under various perturbations in Tabular Question Answering (TQA) tasks.
Exact Match Accuracy (EM): This metric calculates the proportion of instances for which the predicted answer exactly matches the ground truth answer. Formally, if $N$ is the total number of instances, and $\text{correct}(i)$ is an indicator function that is $1$ if the prediction for instance $i$ is correct and 00 otherwise, then:
$\text{EM}=\frac{\sum_{i=1}^{N}\text{correct}(i)}{N}.$
Exact Match Difference (Emd): Let $\text{EM}_{\text{orig}}$ be the EM on the original (unperturbed) dataset, and $\text{EM}_{\text{perturbed}}$ be the EM after applying a perturbation. The Emd quantifies the change in EM due to perturbations:
$\text{Emd}=\text{EM}_{\text{perturbed}}-\text{EM}_{\text{orig}}.$
Negative values indicate performance degradation, while values close to zero imply robustness against perturbations.
Variation Percentage (VP): This metric measures to what extent predictions change after applying perturbations. Given that, C2W is the count of correct before perturbation and wrong after, and W2C is the count of wrong before perturbation and correct after. Given $N$ as the total number of instances, the variation percentage is:
$\text{VP}=\frac{C2W+W2C}{N}.$
A higher VP indicates greater sensitivity of predictions to perturbations, while a lower VP signifies more stable predictions.
Appendix F Performance over Table Complexity
Figure 9: The average EM score for the different structural perturbation over the different size of tables on WTQ, TAT-QA and SCITAB datasets
Figure 10: The average EM scores of the original models and fine-tuned for instruction or conversation models across various table augmentation techniques for WTQ, TAT-QA and SCITAB dataset
Figure 9 presents average model performance across varying table sizes for three tabular question-answering datasets—WTQ, TAT-QA, and SCITAB—under several table structure perturbations. Each plot compares the original condition with different table augmentation operations, such as column swapping, row swapping, and transpose-based modifications.
For WTQ, performance is low overall and highly variable across increasing table sizes, with no single condition consistently outperforming others. In contrast, TAT-QA shows a generally higher and more stable baseline, though still affected by growing table size and perturbations; performance tends to decline as tables grow larger, suggesting sensitivity to complexity. SCITAB results are somewhat mixed, with fluctuations at different size intervals, but the performance remains closer among the different conditions.
Across all three datasets, these results highlight that large language models, though capable, display uneven robustness to structural manipulations of tables. Although instruction tuning and larger model scales improve performance, structural changes continue to pose challenges, indicating the need for more robust, structure-aware approaches for reliable table comprehension.
Appendix G Additional Robustness Result
Figure 11: Comparison of average EM scores across models under varying few-shot settings (0-shot, 1-shot, 2-shot, and 3-shot).
Figure 12: The average cumulative EM score across three different datasets for the different few shot comparisons for various perturbations with three different datasets.
The table 1 summarizes the impact of various perturbation operations on the accuracy and robustness of larger and smaller LLMs models when performing TQA across three distinct datasets (WTQ, TAT-QA, and SCITAB). For each perturbation category, the EM scores typically drop compared to the original configuration, as indicated by negative EmD values. This performance dip is especially notable on the WTQ dataset, where even minor perturbations lead to reduced EM, and certain operations like column transpose and NT conditions cause pronounced declines in EM. Meanwhile, the VP metrics show that predictions shift more substantially under these transformations, signaling sensitivity in both large and small models to altered table structures. Although large models usually have higher baseline EM, they still exhibit performance degradation under perturbations. Particularly striking are conditions like NT, where the table is effectively removed, resulting in EM plummeting to near-zero and high VP, underscoring the models’ heavy reliance on the given tabular data. The results highlight the need for improved architectural and training strategies aimed at enhancing structural awareness and mitigating the negative effects of domain-specific perturbations, thereby making LLMs more robust and reliable in real-world table comprehension scenarios.
Figure 13: The average EM for different models across the TAT-QA dataset with 3-fewshot settings. The models are categorized by their types (Llama2, Llama3, Mistral) and separated by their sizes (7b, 70b, etc.)
Figure 14: Average EM scores for both original and instruction or conversation fine-tuned models on the SCITAB dataset (3-shot setting) under various table augmentation techniques.
G.1 Does newer models have better TQA abilities?
Fig. 15 shows that Llama3 models generally outperform the Llama2 and Mistral models across different configurations, indicating that newer architectures like Llama3 are more effective at table reasoning tasks. The 70b versions of these models generally perform better than their 7b variants, indicating that larger model sizes enhance reasoning capabilities. Overall, the advancements in model architecture and increased model size significantly contribute to better TQA abilities.
Figure 15: The average EM for different models across the WTQ dataset with 3-fewshot settings. The models are categorized by their types (Llama2, Llama3, Mistral) and separated by their sizes (7b, 70b, etc.)
How does model size affect TQA? Larger models (e.g., Llama3-70b, Mixtral-8x7b) generally show higher performance than smaller models (e.g., Llama2-7b, Mistral-7b), as seen in both the bar plot (Fig 15) and heat maps (Fig 10). For instance, Llama3-70b and Mixtral-8x7b have higher EM scores than Llama2-7b and Mistral-7b. This indicates that model size contributes significantly to TQA performance.
Does instruction tuning help TQA? The heat maps in Fig. 10 illustrate the performance of various LLMs, demonstrating that models fine-tuned for instructions or conversation exhibit improved performance. For instance, the Llama3-70b-Instruct model significantly outperforms its original version across all table transformations, indicating that instruction-based fine-tuning enhances the model’s ability to handle complex reasoning tasks. Similarly, conversation-focused fine-tuning also leads to better scores, though with a less pronounced improvement compared to instruction-focused tuning. This suggests that fine-tuning models on specific tasks like following instructions or conversing effectively enhances their capability to interpret and manipulate tabular data, making such approaches valuable for improving performance in structured data tasks. Figure 10(right) distinctly indicates that LLMs that have undergone instruction or conversation-based fine-tuning outperform their base counterparts in TQA tasks for SCITAB dataset. Although SCITAB is inherently a classification task presented in a TQA format, using few-shot prompting provides valuable context, ultimately leading to more accurate and relevant responses.
Appendix H Attention Matrix Measure
Sink rate: Gu et al. introduced the metric that measures the proportion of attention that “collapses" onto a specific token or a small set of tokens. A high sink rate indicates that the attention head focuses very narrowly, potentially ignoring other parts of the input.
$\text{Sink Rate}_{\epsilon}=\frac{1}{T}\sum_{j}\alpha_{1,j}^{(\ell,h)}>\epsilon,$
where $\alpha_{1,j}^{(\ell,h)}$ is the attention weight from the first token (token with index 1(<bos>)) to the j-th token in the $\ell$-th layer and $h$-th attention head.
Entropy: In the context of attention, entropy measures the diversity or distribution of attention across all tokens. High entropy means the attention is distributed more evenly, while low entropy suggests a concentrated attention pattern.
$H_{i}=-\sum_{j}\textbf{A}_{ij}\log(\textbf{A}_{ij})$
where $\textbf{A}_{ij}$ is the attention weight to the $j$-th token.
Frobenius Norm: The Frobenius norm of the attention matrix provides a measure of the overall magnitude or “strength” of attention. It is calculated as the square root of the sum of the squares of all elements in the matrix, indicating the total amount of attention being distributed.
$||\textbf{A}||_{F}=\sqrt{\sum_{i}\sum_{j}\textbf{A}_{ij}^{2}}$
Norm Difference: The difference between the attention matrices computed on the original (dev) and perturbed table(pert):
$\mathrm{NormDiff}\;=\;\bigl{\|}\mathbf{A}^{\text{dev}}\;-\;\mathbf{A}^{\text{pert}}\bigr{\|}_{F},$
where $\mathbf{A}^{dev}$ and $\mathbf{A}^{pert}$ are the attention matrices tokenized using the original table (‘dev”) and the perturbed table, respectively, and $\lVert\cdot\rVert_{F}$ is the Frobenius norm.
H.1 Cross-Model and Dataset Analysis of Perturbation Sensitivity
Figure 16: Spearman correlation between perturbation-induced attention-entropy change and EM degradation across 18 model–dataset pairs.
To evaluate the generality of our findings beyond the Llama3-8B-Instruct model on WTQ, we extend our correlation analysis between attention entropy change and EM performance degradation across a diverse set of model-dataset combinations. These include all the small models defined in 3,for which each models are tested on the WTQ, TATQA, and SCITAB datasets. For each configuration, we compute the Spearman correlation between the perturbation-induced changes in attention entropy and the corresponding drop in EM scores.
As shown in Figure 16, the positive correlation trend persists across nearly all settings. In particular, Instruct and chat variants of the models consistently show the strongest correlation values across datasets, reinforcing their heightened sensitivity to attention dispersion caused by structural perturbations. Notably, even non-instruct variants like Llama2-7B and Mistral-7B exhibit moderate to strong correlations, although the magnitude tends to be lower and more variable. We also find that SCITAB dataset results in most variation in the correlation, mainly due to the performance of the model, with the original table itself having a significantly low performance.
These consistent patterns across models and datasets confirm the robustness of our main claim: perturbation severity that induces greater shifts in attention entropy is reliably associated with a decline in model performance. This suggests that attention entropy change serves as a broadly applicable proxy for evaluating robustness in table-based QA models.
Moreover, this analysis demonstrates that the observed dynamics are not specific to any single dataset or model architecture. Instead, they reflect a more general representational vulnerability within current attention-based architectures when handling perturbed structured inputs.
H.2 Correlation within individual attention
Figure 17: Heatmap showing Spearman correlation between changes in attention entropy and EM difference across all attention heads and layers in the LLama-8B-Instruct model on the WTQ dataset, unsorted to depict actual correlation over each attention head.
Figure 18: Scatter plot showing the correlation between changes in attention entropy and EM scores across five structural perturbations for the WTQ dataset. The plot highlights a strong positive correlation for Layer 0, Head 10, indicating high sensitivity of this head to structural perturbations from 17
Figure 19: Scatter plot showing the correlation between changes in attention entropy and EM scores across five structural perturbations for the WTQ dataset. In contrast to the previous head, Layer 19, Head 23 shows minimal correlation from 17
We find that structural perturbations in tabular inputs affect the internal attention dynamics of the LLMs specifically, Llama-8B-Instruct model on WTQ, and how those dynamics relate to the performance. Figure 17 presents an unsorted Spearman correlation heatmap between per-head changes in attention entropy and the corresponding EM score differences, revealing substantial heterogeneity across all 32 heads and 32 layers. To unpack this variability, Figures 18 and 19 zoom in on the two extremal heads identified by the heatmap: Layer 00, Head $10$ exhibits a strong positive relationship, higher entropy shifts reliably predict larger drops in EM, whereas Layer 19, Head 23 shows significantly less correlation, indicating that its attention distribution remains largely impervious to structural pertubation. Together, these visualizations underscore that certain attention heads are sensitive for model robustness, while others maintain stable focus patterns even under severe perturbations.